EVA Report – November 16th

EVA #9 Report
Today’s EVA would’ve started at 11 AM and ended at 2 PM. Our goal would’ve been to go to Lith Canyon, collect samples from the Salt Wash Member, and return to the Hab. Our EVA for today did not take place, since we needed additional time to clean up the Hab before Survivorman arrives. Also we needed to figure out our transfer arrangements so that we’ll be able to get to Salt Lake City in time to catch their flight. EVA request coming soon!
Prepared by Brandon Ferguson
EVA Officer

GreenHab Report – November 16th

Sol 13 – 16/11/201

Green Hab Officer Atila Meszaros

Ambient: heating 10°C

Ambient with window/door open: door opened at 10:20; closed at 17:34

Shade cloth on/off: on

Average temperature:

Time 10:36 17:06
Temperature inside (°C) 27 15
Temperature outside (°C) 10 8
% Humidity 16 16

Changes to crops:  Crops keep growing!

Daily water usage for crops: 2 galons

Time(s) of watering for crops:  10:36 and 17:06

Changes to research plants: controls keep growing! No signs of germination on the analog soil. The emergency day turned the analog soil like clay, again.

Daily watering and amount of water used: 1 ½ L approximately

Narrative:

Hi, Sylvain! Atila, here. I hope you are having an amazing night.

Kevin’s crops are growing pretty solid, Julio already sent a picture. It’s crazy! I wanted to ask what’s the plan for them.

I thought that the seeds on the Petri dishes were ready to be planted on the analog soil, again. They just needed another one more day. The quinoa seeds are not germinating on the control, either. I’m guessing that the day I didn’t water them was enough for them from drying. That almost killed all the project! I will need to continue with my experiment on my university with the analog soil to get solid results.

I think that’s pretty much all, if you need more information about the GH I will be glad to send it to you.

Thank you for being there, Sylvain!

Have a beautiful night on Earth.

Support/supplies needed: none

Operations Report – November 16th

Crew 182 Operations Report  16Nov2017

SOL: 13

Name of person filing report:  Carmen Atauconcha

Non-nominal systems:

WATER HEATER: Finally, we found the problem. The new water heater is not installed, yet.

CHARGERS OF THE SUITS: Two of the chargers are working, but we moved them and now it is hard for us to connect them, again.

 

Generator—  It worked the whole night since 16:30pm of 15Nov2017 until 12:00 pm of 16Nov2017.  Turn on at 16:50 pm 16Nov2017.  Operand.

Solar—

READING FROM 16 NOV 2017

 

TIME SOC READINGS
12:00 pm 100%
13:00 pm 98%
14:00 pm 91%
15:00 pm 84%
16:20 pm 80%
16:50 pm 75%

 

Diesel – 50 %

Propane – 49 psi

Gasoline (5 Gallon containers for ATV) –  2 ½ gallons.

Water (trailer) –  0 gallons

Water (static) – 300  gallons

Trailer to Static Pump used –   No

Water (loft) – Static to Loft Pump used –  yes

Water Meter:  60 gallons

Toilet tank emptied:  yes

ATV’s Used: No

Oil Checked:  No

Atv # Fuel Used Gals:  0  gallons

Tires Status: Perfect

# Hours the ATVs were Used today: 0 hours

Deimos used:  No

Hours: 0

Beginning charge:  –

Ending charge:  –

Currently charging: Yes
HabCar used? No

Notes and Comments:

I would call today, the day of cleaning. Everyone here is tired. Literally, we cleaned every hole here. There was too much dust. Today in the morning, the commander went to search transportation to return to the Earth.  We are sad to leave Mars, and at the same time, happy for returning at home.

From another hand, survivor man is coming. We are so excited¡

 

Summary of Hab Operations:
We cleaned the whole hab. We found, a couple of glasses and gloves, behind the area of the tools.

Summary of GreenHab operations:

We cleaned the greenhouse. We found too much dust. The two wheelbarrows are in the greenhouse, now.

Summary of ScienceDome operations:

The science dome is finally cleaned. In fact, we found tons of dust under the carpet.

We are using the lamps of the Hab, for the laboratory.

Summary of health and safety issues:

Now, we can say the station is in equilibrium with our health and safety.

Questions, concerns and requests to Mission Support:

We need gasoline for the ATV’S, please. URGENT

Sol Summary – November 16th

MDRS Daily Summary Report for Sol 13.

Summary Title: Organizing MDRS.

Mission Status: Preparing to leave MDRS. We did some improvements in MDRS fixing some issues in the physical facilities.

Sol Activity Summary: We just concluded the cleaning procedures of the entire installation, also cleaning the GreenHab and the laboratory. We developed many different activities organizing and fixing the facilities of MDRS.

Look Ahead Plan: Tomorrow, probably will happen an EVA, but We are just waiting the SurvivorMan to decide about would be interesting to him.

Anomalies in work: Nope.

Weather: 17o C, 62o F.

Crew Physical Status: Fine.

EVA: We planed an EVA today, but not was possible. We planed to go to Lith Canyon, collect samples from the Salt Wash Member, and return to the Hab. Our EVA for today did not take place, since we needed additional time to clean up the Hab before Survivorman arrives.

 

Reports to be file:

Commanders Report

Daily Summary Report

Operation Report

GreenHab Report

HSO Report

XO Report

Crew-photos Crew 182 16Nov2017.

 

Support Requested: Nope.